8/24 equals 14/x
What is x and how did you get the answer?

Answer by stanbon(75887)

8/24 equals 14/x
What is x and how did you get the answer?

Invert both sides to get:
x/14 = 8/24
------
Multiply both sides by 14 to get:
x = 14(8/24)
------
x = 14(1/3)
----
x = 4 2/3
-------------
Cheers,
Stan H.
----------